OpenStudy (anonymous):

How many solutions does the equation have? 6-a=1/3(9a-18) Please no bs . I'm having a bit of a panic attack here so if you could just answer as simply as you can , that would be positively wonderful

OpenStudy (popitree):

this is a quadratic equation-- max power of a = 2, so 2 solution

OpenStudy (anonymous):

I don't need the solution , just the possible amount of solutions ..

OpenStudy (shinalcantara):

popitree is right.there are 2 possible values of a since it's quadratic

OpenStudy (popitree):

2 is not solution.. that is no of solutions

OpenStudy (popitree):

:)

OpenStudy (anonymous):

So the answer is no solution ?

OpenStudy (popitree):

the equation has two solutions

OpenStudy (anonymous):

The only possible answers are no solution , infinite solutions , and one solution ..

jhonyy9 (jhonyy9):

6-a = (9a-18)/3 6-a = 3(3a-6)/3 simplified 3 by 3 then will get 6-a = 3a-6 12 = 4a a= 12/4 a = 3 hope this is now sure understandably